OpenStudy (anonymous):

I Will Give Medal!!!! Tiffany kicks a soccer ball off the ground and in the air with an initial velocity of 28 feet per second. Using the formula H(t)=-16^2+vt+s, what is the maximum height the soccer ball reaches? A. 15.6 feet B. 14.1 feet C. 13.5 feet D. 12.3 feet

OpenStudy (jlg030597):

Whats your guess buddy

OpenStudy (jlg030597):

Ill help you get to the answer if your wrong

OpenStudy (anonymous):

Honestly I forget how to use the formula.

OpenStudy (mathmale):

Are you sure it's H(t)=-16^2+vt+s? do you want that "v" in there, or was that a typo?

OpenStudy (anonymous):

yes the v is in the problem

OpenStudy (mathmale):

And that's an "s" ? With two unknowns, it's pretty difficult to estimate what the max height would be. If you had H(x) = -16t^2 + 28t + 5 (the 5 arbitrary), you could graph H(x) and easily find the max height. But without knowing the value of " s ", you can't. Double check, please, that you're copied the problem down correctly.

OpenStudy (mathmale):

Is this course "algebra" or "calculus"?

OpenStudy (anonymous):

I think they put the s there so that it would get replaced like by that 5. this is algebra 1

OpenStudy (mathmale):

that "s" represents the INITIAL HEIGHT of the soccer ball. Obviously it's not going to be a very large number, since most of us could not kick a soccer ball that was more than about 5 feet in the air, right? If v = initial velocity = 28 (ft/sec), then your equation becomes APPROXIMATELY H(t) = -16*t^2+28t+5. Note that you didn't include the t^2 earlier, so I've fixed that. Now, if you'd graph H(t) = -16*t^2+28t+5, you could see pretty easily what the maximum height will be. Sound OK?

OpenStudy (anonymous):

yeah ok

OpenStudy (mathmale):

Are you now able to estimate the max. ht.?

OpenStudy (mathmale):

I've graphed this function on my TI-83 Plus calculator and see that the graph is a parabola that opens down. Using the Trace function, it's easy to determine the approx. max. height reached by the soccer ball. For s = 5 feet, I get max. ht. = approx. 17 feet. That's reasonable for a soccer ball kick, isn't it? Maybe you can kick a soccer ball 100 ft into the air; sadly, I can't.

OpenStudy (anonymous):

Haha alright so is that it is 17 the anwser

OpenStudy (mathmale):

Won't say! You'll have to decide that for yourself!

OpenStudy (anonymous):

Well I guess not being that its not any of the choices haha

OpenStudy (mathmale):

Choose the answer choice that makes the most sense to you, even if it's not exactly 17 ft!